OpenStudy (freemap):

Which two points are on the graph of M(a)=-3/4-12? Choose exactly two answers that are correct. A. (–16, 0) B. (–9, 0) C. (0, 12) D. (0, –12)

OpenStudy (callisto):

What is the slope of the line in the given equation?

OpenStudy (freemap):

I don't have two points so would I substitute 0?

OpenStudy (callisto):

Nope, from the given equation: M(a) = (-3/4)a - 12

OpenStudy (freemap):

I just don't understand how you find slope with that like, you're not doing like y2-y1 then x2-x1. I don't understand how to solve this

OpenStudy (texaschic101):

M(a) is the same as y...so your equation would be y = -3/4a - 12 and in y = mx + b form,although in this problem it is y = ma + b, the slope is in the m position.

OpenStudy (texaschic101):

wait.....I think you can just sub in your points... y = -3/4(a stands for x) + b now sub in your points and see which is true

OpenStudy (freemap):

ok, so like -3/4(-16)-0?

OpenStudy (texaschic101):

no.. like this... y = -3/4x - 12 (-16,0)...x = -16 and y = 0 0 = -3/4(-16) - 12 0 = 12 - 12 0 = 0 (correct)

OpenStudy (freemap):

ok I'll attempt the next one just a sec

OpenStudy (freemap):

Ok so no because 0=6.75-12

OpenStudy (texaschic101):

correct...so that is not a solution

OpenStudy (texaschic101):

you get it now ? Can you finish ?

OpenStudy (freemap):

Yes I can finish Thanks for the help

OpenStudy (texaschic101):

just remember, in a function, whether it is f(x) or m(a) or anything like that, it represents y

OpenStudy (freemap):

Thank You

OpenStudy (texaschic101):

sure thing :)